# STAR-RIS Enhanced UAV-Enabled MEC Networks with Bi-Directional Task Offloading

### v1

@@ III-B2 Designing $\boldsymbol{\Xi}_{2}$ with given $\boldsymbol{\zeta}$ and $\mathbf{Q}$
$\max_{\boldsymbol{\Xi}_{2}}~{}~{}\frac{L_{\mathrm{tol}}(\boldsymbol{\Xi}_{2})}{E_{\mathrm{tol}}(\boldsymbol{\Xi}_{2})}$
@@
Note that, problem (III-B2) is a non-convex optimization problem because of the non-convexity of the fractional objective function and constraints (14), (15) and (26f). However, we can derive the close form expression of the optimal reflected and transmitted phases according to the following Theorem 1.
@@
The obtained optimal reflection and transmission phases at the $n$-th time slot can be given by
$\boldsymbol{\phi}_{\mathrm{r}}[n]=$ $\operatorname{arg}(\operatorname{norm}(\mathbf{h}_{\mathrm{rb}}^{*}[n]\circ\mathbf{h}_{\mathrm{r}k}[n])),$ (33)
$\boldsymbol{\phi}_{\mathrm{t}}[n]=$ $\operatorname{arg}(\operatorname{norm}(\mathbf{h}_{\mathrm{ru}}^{*}\circ\mathbf{h}_{\mathrm{r}k}[n])),$ (34)
@@
To maximize the achievable offloaded rates $R^{\mathrm{ua}}_{k}[n]$ and $R^{\mathrm{BS}}_{k}[n]$ when $\zeta_{k}[n]=1$ in the $n$-th time slot, the optimal phases of the STAR-RIS’s passive beamforming can be determined by utilizing the maximum ratio transmission (MRT) technique. ∎
Algorithm 1: The Proposed Iterative Algorithm for Solving the Sub-problem (III-B1)
1: Initialize the feasible point $\big{(}\mathbf{L}^{(0,0)},\boldsymbol{\zeta}^{(0,0)}\big{)}$; Define the tolerance accuracy thresholds as $\varepsilon$ and $\widehat{\varepsilon}$; Set the outer iteration index $l$ = 0.
2: While $v_{1}>\varepsilon_{1}$ or $l=0$ do
3:  Set the inner iteration index $t$ = 0; Initialize $\hat{\rho}^{(0)}$.
4:  While $v_{2}>\varepsilon_{2}$ or $t=0$ do
5:   Solve the optimization problem (III-B1) with the given
     $\big{(}\mathbf{L}_{\mathrm{b}}^{(l,t)},\boldsymbol{\zeta}^{(l,t)}\big{)}$ and update $\big{(}\mathbf{L}_{\mathrm{b}}^{(l,t+1)},\boldsymbol{\zeta}^{(l,t+1)}\big{)}$ with
     the obtained solutions.
6:   Calculate $v_{2}=\max\limits_{k\in\mathcal{K},n\in\mathcal{N}}\eta_{k}[n]$ based on the acquired
     Let $t=t+1$.
7:  end while
8:  Update $\big{(}\mathbf{L}_{\mathrm{b}}^{(l+1,0)},\boldsymbol{\zeta}^{(l+1,0)}\big{)}$ with $\big{(}\mathbf{L}_{\mathrm{b}}^{(l,t)},\boldsymbol{\zeta}^{(l,t)}\big{)}$.
9:  Calculate $v_{1}=L_{\mathrm{tol}}^{(l+1)}-\psi^{(l)}E_{\mathrm{tol}}^{(l+1)}$ based on the acq-
  uired solution;
10: Calculate the $\psi^{(l+1)}=\frac{L_{\mathrm{tol}}(\mathbf{L}_{\mathrm{b}}^{(l+1,0)},\boldsymbol{\zeta}^{(l+1,0)})}{E_{\mathrm{tol}}(\mathbf{L}_{\mathrm{b}}^{(l+1,0)},\boldsymbol{\zeta}^{(l+1,0)})}$; Let $l=$
  $l+1$.
12: end while
Next, we will focus on designing the reflection and transmission amplitudes of STAR-RIS’s passive beamforming, i.e., $\{\beta^{m}_{\mathrm{\kappa}}[n]\}_{m\in\mathcal{M},n\in\mathcal{N}}$. In terms of the fractional objective function, similarly, we resort to the Dinkelbach’s algorithm to transform it. For the non-convex constraints (14) and (15), we first rewrite $\boldsymbol{\vartheta}_{\kappa}[n]=\boldsymbol{\Phi}_{\kappa}[n]\boldsymbol{\beta}_{\kappa}[n],\kappa\in\{\mathrm{r},\mathrm{t}\}$, where we have $\boldsymbol{\vartheta}_{\kappa}[n]=\operatorname{diag}(\boldsymbol{\Theta}_{\kappa}[n])$, $\boldsymbol{\beta}_{\kappa}[n]=\{\beta_{\kappa}^{1}[n],\cdots,\beta_{\kappa}^{m}[n],\cdots,\beta_{\kappa}^{M}[n]\}^{T}$, and $\boldsymbol{\Phi}_{\kappa}[n]=\operatorname{Diag}(e^{j\boldsymbol{\phi}_{\kappa}[n]})$. Therefore, $R^{\mathrm{BS}}_{k}[n]$ and $R^{\mathrm{ua}}_{k}[n]$ can be further re-expressed as
@@
$R^{\mathrm{ua}}_{k}[n]=$ $\zeta_{k}[n]B\log_{2}\Big{(}1+\boldsymbol{\beta}_{\mathrm{t}}^{T}[n]\mathbf{E}_{k}[n]\boldsymbol{\beta}_{\mathrm{t}}[n]\Big{)},$ (36)
@@
  • $\mathbf{E}_{k}[n]=\frac{\boldsymbol{\Phi}_{\mathrm{t}}^{H}[n]}{\sigma_{\mathrm{ua}}^{2}}(\mathbf{h}_{\mathrm{ru}}^{*}\circ\mathbf{h}_{\mathrm{r}k}[n])(\mathbf{h}_{\mathrm{ru}}^{*}\circ\mathbf{h}_{\mathrm{r}_{k}}[n])^{H}\boldsymbol{\Phi}_{\mathrm{t}}[n]$.
Then, we introduce auxiliary variables $\gamma^{\mathrm{BS}}_{k}[n]$ and $\gamma^{\mathrm{ua}}_{k}[n]$ which satisfy $\gamma^{\mathrm{BS}}_{k}[n]\leq\boldsymbol{\beta}_{\mathrm{r}}^{T}[n]\mathbf{F}_{k}[n]\boldsymbol{\beta}_{\mathrm{r}}[n]$ and $\gamma^{\mathrm{ua}}_{k}[n]\leq\boldsymbol{\beta}_{\mathrm{t}}^{T}[n]\mathbf{E}_{k}[n]\boldsymbol{\beta}_{\mathrm{t}}[n]$. Thus, the problem designing reflection and transmission amplitudes in the ($\hat{t}+1$)-th iteration of the Dinkelbach’s algorithm can be equivalently transformed as
$~{}\max_{\boldsymbol{\beta}_{\mathrm{r}},\boldsymbol{\beta}_{\mathrm{t}},\boldsymbol{\gamma}}L_{\mathrm{tol}}(\boldsymbol{\beta}_{\mathrm{r}},\boldsymbol{\beta}_{\mathrm{t}},\boldsymbol{\gamma})-\hat{\psi}^{(\hat{t})}E_{\mathrm{tol}}(\boldsymbol{\beta}_{\mathrm{r}},\boldsymbol{\beta}_{\mathrm{t}},\boldsymbol{\gamma})$
@@
$~{}~{}\gamma^{\mathrm{ua}}_{k}[n]\leq\boldsymbol{\beta}_{\mathrm{t}}^{T}[n]\mathbf{E}_{k}[n]\boldsymbol{\beta}_{\mathrm{t}}[n],~{}\forall k\in\mathcal{K},n\in\mathcal{N},$ (37d)
$~{}~{}\gamma^{\mathrm{BS}}_{k}[n]\leq\boldsymbol{\beta}_{\mathrm{r}}^{T}[n]\mathbf{F}_{k}[n]\boldsymbol{\beta}_{\mathrm{r}}[n],~{}\forall k\in\mathcal{K},n\in\mathcal{N},$ (37e)
$~{}~{}(\beta_{\mathrm{r}}^{m}[n])^{2}+(\beta_{\mathrm{t}}^{m}[n])^{2}\leq 1,~{}\forall m\in\mathcal{M},n\in\mathcal{N}.$ (37f)
where $\boldsymbol{\gamma}\triangleq\{\gamma^{\mathrm{ua}}_{k}[n],\gamma^{\mathrm{BS}}_{k}[n],k\in\mathcal{K},n\in\mathcal{N}\}$. Actually, problem (III-B2) is still a non-convex optimization problem due to the non-convex constraints (37d) and (37e). In order to address this issue, we employ a linear lower bound, specifically the first-order Taylor expansion, to approximate the right-hand side of constraints (37d) and (37e) and replace them accordingly. Note that the equality sign in constraint (26f) is replaced with an inequality sign to form constraint (37f). In fact, this change does not impact the fulfilment of the equality constraint, because the constraint (37f) is satisfied with strict equality in the optimal solution of problem (III-B2). Consequently, we can obtain a convex optimization problem.
@@ III-B3 Designing $\boldsymbol{\Xi}_{3}$ with given $\boldsymbol{\zeta}$ and $\boldsymbol{\Upsilon}$
$\max_{\boldsymbol{\Xi}_{3}}~{}~{}~{}\frac{L_{\mathrm{tol}}(\boldsymbol{\Xi}_{3})}{E_{\mathrm{tol}}(\boldsymbol{\Xi}_{3})},$
@@
In fact, problem (III-B3) is non-convex because of the objective function and constraints (14) and (15). Similarly, we utilize the Dinkelbach’s algorithm to transform the fractional objective function. In order to handle this non-convex constraints, we first introduce the auxiliary variables $\lambda_{k}[n]$ and $\tilde{\lambda}[n]$ with
@@
Actually, $\tilde{R}_{k}^{\mathrm{ua}}[n]$ is a convex function w.r.t. $\lambda_{k}[n]$ and $\tilde{R}_{k}^{\mathrm{BS}}[n]$ is the jointly convex function w.r.t. $\lambda_{k}[n]$ and $\tilde{\lambda}[n]$, and thus we can apply the first-order Taylor expansion of $\tilde{R}_{k}^{\mathrm{ua}}[n]$ and $\tilde{R}_{k}^{\mathrm{BS}}[n]$ at the given point $(\lambda_{k}^{(s)}[n],\tilde{\lambda}^{(s)}[n])$ in $(s+1)$-th iteration of the Dinkelbach’s algorithm to convert constraints (14) and (15) as the following convex constraints
$\zeta_{k}[n]\delta_{\mathrm{t}}\widehat{R}_{k}^{\mathrm{ua}}[n]\geq l_{k}^{\mathrm{ua}}[n],~{}\forall k\in\mathcal{K},n\in\mathcal{N},$ (43)
$\zeta_{k}[n]\delta_{\mathrm{t}}\widehat{R}_{k}^{\mathrm{BS}}[n]\geq l_{k}^{\mathrm{BS}}[n],~{}\forall k\in\mathcal{K},n\in\mathcal{N},$ (44)
@@
  • $\widehat{R}_{k}^{\mathrm{ua}}[n]=\log_{2}\Big{(}1+\frac{\rho p\left|\mathbf{h}^{H}_{\mathrm{ru}}\boldsymbol{\Theta}_{\mathrm{t}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}}{\lambda^{(s)}_{k}[n]\sigma_{\mathrm{ua}}^{2}}\Big{)}+(\lambda_{k}[n]-\lambda^{(s)}_{k}[n])f(\lambda^{(s)}_{k}[n])$.
  • $\widehat{R}_{k}^{\mathrm{BS}}[n]=(\lambda_{k}[n]-\lambda^{(s)}_{k}[n])\tilde{f}_{1}(\lambda^{(s)}_{k}[n],\tilde{\lambda}^{(s)}[n])+(\tilde{\lambda}[n]-\tilde{\lambda}^{(s)}[n])\tilde{f}_{2}(\lambda^{(s)}_{k}[n],\tilde{\lambda}^{(s)}[n])+\log_{2}\big{(}1+\frac{\rho^{2}p\left|\widehat{\mathbf{h}}_{\mathrm{BR}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}}{\lambda^{(s)}_{k}[n]\tilde{\lambda}^{(s)}[n]\sigma_{\mathrm{ua}}^{2}}\big{)}$.
  • $f(\lambda^{(s)}_{k}[n])=\frac{-\rho p\left|\mathbf{h}^{H}_{\mathrm{ru}}\boldsymbol{\Theta}_{\mathrm{t}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}}{\ln 2\big{(}\rho p\left|\mathbf{h}^{H}_{\mathrm{ru}}\boldsymbol{\Theta}_{\mathrm{t}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}+\lambda^{(s)}_{k}[n]\sigma_{\mathrm{ua}}^{2}\big{)}\lambda^{(s)}_{k}[n]}$.
  • $\tilde{f}_{1}(\lambda^{(s)}_{k}[n],\tilde{\lambda}^{(s)}[n])=\\ \frac{-\rho^{2}p\left|\widehat{\mathbf{h}}_{\mathrm{BR}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}}{\ln 2\lambda^{(s)}_{k}[n]\big{(}\rho^{2}p\left|\widehat{\mathbf{h}}_{\mathrm{BR}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}+\lambda^{(s)}_{k}[n]\tilde{\lambda}^{(s)}[n]\sigma_{\mathrm{ua}}^{2}\big{)}}$.
  • $\tilde{f}_{2}(\lambda^{(s)}_{k}[n],\tilde{\lambda}^{(s)}[n])=\\ \frac{-\rho^{2}p\left|\widehat{\mathbf{h}}_{\mathrm{BR}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}}{\ln 2\tilde{\lambda}^{(s)}[n]\big{(}\rho^{2}p\left|\widehat{\mathbf{h}}_{\mathrm{BR}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}+\lambda^{(s)}_{k}[n]\tilde{\lambda}^{(s)}[n]\sigma_{\mathrm{ua}}^{2}\big{)}}$.
Note that, $E_{\mathrm{tol}}$ is a non-convex function w.r.t. variable $\mathbf{Q}$ due to the non-convexity of $E_{\mathrm{ua}}^{\mathrm{fly}}$. To tackle this problem, we futher introduce a non-negative auxiliary variable $\widehat{\mu}[n]$ with $\widehat{\mu}^{2}[n]\geq(1+\frac{v^{4}[n]}{4v_{0}^{4}})^{\frac{1}{2}}-\frac{v^{2}[n]}{2v_{0}^{2}}$ to obtain the upper bound of $E_{\mathrm{ua}}^{\mathrm{fly}}$, which is expressed as
@@
$g(\widehat{\mu},v[n])=(\widehat{\mu}^{(s)}[n])^{2}+2\mu^{(s)}[n](\widehat{\mu}[n]-\widehat{\mu}^{(s)}[n])$
$+\frac{2}{v_{0}^{2}\delta_{\mathrm{t}}^{2}}(\mathbf{q}_{\mathrm{ua}}^{(s)}[n+1]-\mathbf{q}_{\mathrm{ua}}^{(s)}[n])^{T}(\mathbf{q}_{\mathrm{ua}}[n+1]-\mathbf{q}_{\mathrm{ua}}[n])$
$-\frac{\left\|\mathbf{q}_{\mathrm{ua}}^{(s)}[n+1]-\mathbf{q}_{\mathrm{ua}}^{(s)}[n]\right\|^{2}}{v_{0}^{2}\delta_{\mathrm{t}}^{2}}\geq\frac{1}{\widehat{\mu}^{2}[n]}.$ (46)
As a result, the optimization problem (III-B3) in the $(s+1)$-th iteration of the Dinkelbach’s algorithm can be transformed as
$\max_{\boldsymbol{\Xi}}~{}~{}~{}\mathbf{L}(\boldsymbol{\Xi})-\tau^{(s+1)}\widehat{E}_{\mathrm{tol}}(\boldsymbol{\Xi}),$

### v3

@@ III-B2 Designing $\boldsymbol{\Xi}_{2}$ with given $\boldsymbol{\zeta}$ and $\mathbf{Q}$
$\max_{\boldsymbol{\Xi}_{2}}~{}~{}L_{\mathrm{tol}}(\boldsymbol{\Xi}_{2})-\psi^{(l)}E_{\mathrm{tol}}(\boldsymbol{\Xi}_{2}),$
@@
Note that, problem (34) is a non-convex optimization problem because of the non-convexity of constraints (14), (15) and (26f). However, we can derive the close-form expression of the optimal reflected phases according to the Theorem 1.
@@
The obtained optimal reflection phases at the $n$-th time slot can be given by
$\boldsymbol{\phi}_{\mathrm{r}}^{\mathrm{opt}}[n]=$ $\operatorname{arg}(\operatorname{norm}(\mathbf{h}_{\mathrm{rb}}^{*}[n]\circ\mathbf{h}_{\mathrm{r}k}[n])),$ (35)
@@
The proof of Theorem 1 is given by Appendix A
Next, we will focus on designing the reflection amplitudes and transmitted coefficient of STAR-RIS’s passive beamforming, i.e., $\{\beta^{m}_{\mathrm{r}}[n]\}_{m\in\mathcal{M},n\in\mathcal{N}}$ and $\{\boldsymbol{\Theta}_{\mathrm{t}}[n]\}_{n\in\mathcal{N}}$. For the non-convex constraints (14) and (15), we first rewrite $\boldsymbol{\vartheta}_{\mathrm{r}}[n]=\boldsymbol{\Phi}_{\mathrm{r}}[n]\boldsymbol{\beta}_{\mathrm{r}}[n]$, where $\boldsymbol{\vartheta}_{\mathrm{r}}[n]=\operatorname{diag}(\boldsymbol{\Theta}_{\mathrm{r}}[n])$, $\boldsymbol{\beta}_{\mathrm{r}}[n]=[\beta_{\mathrm{r}}^{1}[n],\cdots,\beta_{\mathrm{r}}^{m}[n],\cdots,\beta_{\mathrm{r}}^{M}[n]]^{T}$, and $\boldsymbol{\Phi}_{\mathrm{r}}[n]=\operatorname{Diag}(e^{j\boldsymbol{\phi}_{\mathrm{r}}^{\mathrm{opt}}[n]})$. Therefore, $R^{\mathrm{BS}}_{k}[n]$ and $R^{\mathrm{ua}}_{k}[n]$ can be further re-expressed as
@@
$R^{\mathrm{ua}}_{k}[n]=$ $\zeta_{k}[n]B\log_{2}\Big{(}1+\boldsymbol{\vartheta}_{\mathrm{t}}^{H}[n]\mathbf{E}_{k}[n]\boldsymbol{\vartheta}_{\mathrm{t}}[n]\Big{)},$ (37)
@@
  • $\boldsymbol{\vartheta}_{\mathrm{t}}[n]=\operatorname{diag}(\boldsymbol{\Theta}_{\mathrm{t}}[n])=\{\beta^{1}_{\mathrm{t}}[n]e^{j\phi^{1}_{\mathrm{t}}[n]},\cdots,\\ \beta^{m}_{\mathrm{t}}[n]e^{j\phi^{m}_{\mathrm{t}}[n]},\cdots,\beta^{M}_{\mathrm{t}}[n]e^{j\phi^{M}_{\mathrm{t}}[n]}\}$,
@@
  • $\mathbf{E}_{k}[n]=(\mathbf{h}_{\mathrm{ru}}^{*}\circ\mathbf{h}_{\mathrm{r}k}[n])(\mathbf{h}_{\mathrm{ru}}^{*}\circ\mathbf{h}_{\mathrm{r}_{k}}[n])^{H}$.
Then, we introduce auxiliary variables $\gamma^{\mathrm{BS}}_{k}[n]$ and $\gamma^{\mathrm{ua}}_{k}[n]$ which satisfy $\gamma^{\mathrm{BS}}_{k}[n]\leq\boldsymbol{\beta}_{\mathrm{r}}^{T}[n]\mathbf{F}_{k}[n]\boldsymbol{\beta}_{\mathrm{r}}[n]$ and $\gamma^{\mathrm{ua}}_{k}[n]\leq\boldsymbol{\vartheta}_{\mathrm{t}}^{H}[n]\mathbf{E}_{k}[n]\boldsymbol{\vartheta}_{\mathrm{t}}[n]$. Thus, the problem designing reflection amplitudes and transmission coefficients in the ($l+1$)-th iteration of the Dinkelbach’s algorithm can be equivalently transformed as
$~{}\max_{\mathbf{L},\boldsymbol{\beta}_{\mathrm{r}},\boldsymbol{\vartheta}_{\mathrm{t}},\boldsymbol{\gamma}}L_{\mathrm{tol}}(\boldsymbol{\beta}_{\mathrm{r}},\boldsymbol{\vartheta}_{\mathrm{t}},\boldsymbol{\gamma})-\psi^{(l)}E_{\mathrm{tol}}(\boldsymbol{\beta}_{\mathrm{r}},\boldsymbol{\vartheta}_{\mathrm{t}},\boldsymbol{\gamma})$
@@
$~{}~{}\gamma^{\mathrm{ua}}_{k}[n]\leq\boldsymbol{\vartheta}_{\mathrm{t}}^{H}[n]\mathbf{E}_{k}[n]\boldsymbol{\vartheta}_{\mathrm{t}},~{}\forall k\in\mathcal{K},~{}n\in\mathcal{N},$ (38d)
$~{}~{}\gamma^{\mathrm{BS}}_{k}[n]\leq\boldsymbol{\beta}_{\mathrm{r}}^{T}[n]\mathbf{F}_{k}[n]\boldsymbol{\beta}_{\mathrm{r}}[n],~{}\forall k\in\mathcal{K},~{}n\in\mathcal{N},$ (38e)
$~{}~{}(\beta_{\mathrm{r}}^{m}[n])^{2}+(\beta_{\mathrm{t}}^{m}[n])^{2}\leq 1,~{}\forall m\in\mathcal{M},~{}n\in\mathcal{N}.$ (38f)
where $\boldsymbol{\gamma}\triangleq\{\gamma^{\mathrm{ua}}_{k}[n],\gamma^{\mathrm{BS}}_{k}[n],k\in\mathcal{K},n\in\mathcal{N}\}$. Actually, problem (38) is still a non-convex optimization problem due to the non-convex constraints (38d) and (38e). In order to address this issue, we employ a linear lower bound, specifically the first-order Taylor expansion, to approximate the right-hand side of constraints (38d) and (38e) and replace them accordingly. It is worth noting that the equality sign in constraint (26f) has been substituted with an inequality sign to establish the convex constraint (38f), transforming (38) into a convex optimization problem.
      Proposition 1.
In fact, this replacement of the equal sign does not impact the fulfilment of the equality constraint, because the constraint (38f) is satisfied with strict equality in the optimal solution of problem (38).
Proof.
The proof of Proposition 1 is given by Appendix B. ∎
@@ III-B3 Designing $\boldsymbol{\Xi}_{3}$ with given $\boldsymbol{\zeta}$ and $\boldsymbol{\Upsilon}$
$\max_{\boldsymbol{\Xi}_{3}}~{}L_{\mathrm{tol}}(\boldsymbol{\Xi}_{3})-\psi^{(l)}E_{\mathrm{tol}}(\boldsymbol{\Xi}_{3}),,$
@@
In fact, problem (39) is non-convex because of constraints (14) and (15). In order to handle this non-convex constraints, we first introduce the auxiliary variables $\lambda_{k}[n]$ and $\tilde{\lambda}[n]$ with
@@
Actually, $\tilde{R}_{k}^{\mathrm{ua}}[n]$ is a convex function w.r.t. $\lambda_{k}[n]$ and $\tilde{R}_{k}^{\mathrm{BS}}[n]$ is the jointly convex function w.r.t. $\lambda_{k}[n]$ and $\tilde{\lambda}[n]$, and thus we can apply the first-order Taylor expansion of $\tilde{R}_{k}^{\mathrm{ua}}[n]$ and $\tilde{R}_{k}^{\mathrm{BS}}[n]$ at the given point $(\lambda_{k}^{(l)}[n],\tilde{\lambda}^{(l)}[n])$ in $(l+1)$-th iteration of the Dinkelbach’s algorithm to convert constraints (14) and (15) as the following convex constraints
$\zeta_{k}[n]\delta_{\mathrm{t}}\widehat{R}_{k}^{\mathrm{ua}}[n]\geq l_{k}^{\mathrm{ua}}[n],~{}\forall k\in\mathcal{K},~{}n\in\mathcal{N},$ (44)
$\zeta_{k}[n]\delta_{\mathrm{t}}\widehat{R}_{k}^{\mathrm{BS}}[n]\geq l_{k}^{\mathrm{BS}}[n],~{}\forall k\in\mathcal{K},~{}n\in\mathcal{N},$ (45)
@@
  • $\widehat{R}_{k}^{\mathrm{ua}}[n]=\log_{2}\Big{(}1+\frac{\rho p\left|\mathbf{h}^{H}_{\mathrm{ru}}\boldsymbol{\Theta}_{\mathrm{t}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}}{\lambda^{(l)}_{k}[n]\sigma_{\mathrm{ua}}^{2}}\Big{)}+(\lambda_{k}[n]-\lambda^{(s)}_{k}[n])f(\lambda^{(l)}_{k}[n])$.
  • $\widehat{R}_{k}^{\mathrm{BS}}[n]=(\lambda_{k}[n]-\lambda^{(l)}_{k}[n])\tilde{f}_{1}(\lambda^{(l)}_{k}[n],\tilde{\lambda}^{(l)}[n])+(\tilde{\lambda}[n]-\tilde{\lambda}^{(l)}[n])\tilde{f}_{2}(\lambda^{(l)}_{k}[n],\tilde{\lambda}^{(l)}[n])+\log_{2}\big{(}1+\frac{\rho^{2}p\left|\widehat{\mathbf{h}}_{\mathrm{BR}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}}{\lambda^{(l)}_{k}[n]\tilde{\lambda}^{(l)}[n]\sigma_{\mathrm{ua}}^{2}}\big{)}$.
  • $f(\lambda^{(l)}_{k}[n])=\frac{-\rho p\left|\mathbf{h}^{H}_{\mathrm{ru}}\boldsymbol{\Theta}_{\mathrm{t}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}}{\ln 2\big{(}\rho p\left|\mathbf{h}^{H}_{\mathrm{ru}}\boldsymbol{\Theta}_{\mathrm{t}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}+\lambda^{(l)}_{k}[n]\sigma_{\mathrm{ua}}^{2}\big{)}\lambda^{(l)}_{k}[n]}$.
  • $\tilde{f}_{1}(\lambda^{(l)}_{k}[n],\tilde{\lambda}^{(l)}[n])=\\ \frac{-\rho^{2}p\left|\widehat{\mathbf{h}}_{\mathrm{BR}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}}{\ln 2\lambda^{(l)}_{k}[n]\big{(}\rho^{2}p\left|\widehat{\mathbf{h}}_{\mathrm{BR}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}+\lambda^{(l)}_{k}[n]\tilde{\lambda}^{(l)}[n]\sigma_{\mathrm{ua}}^{2}\big{)}}$.
  • $\tilde{f}_{2}(\lambda^{(l)}_{k}[n],\tilde{\lambda}^{(l)}[n])=\\ \frac{-\rho^{2}p\left|\widehat{\mathbf{h}}_{\mathrm{BR}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}}{\ln 2\tilde{\lambda}^{(l)}[n]\big{(}\rho^{2}p\left|\widehat{\mathbf{h}}_{\mathrm{BR}}^{H}[n]\boldsymbol{\Theta}_{\mathrm{r}}^{H}[n]\widehat{\mathbf{h}}_{\mathrm{r}k}[n]\right|^{2}+\lambda^{(l)}_{k}[n]\tilde{\lambda}^{(l)}[n]\sigma_{\mathrm{ua}}^{2}\big{)}}$.
Note that, $E_{\mathrm{tol}}$ is a non-convex function w.r.t. variable $\mathbf{Q}$ due to the non-convexity of $E_{\mathrm{ua}}^{\mathrm{fly}}$. To tackle this problem, we further introduce a non-negative auxiliary variable $\widehat{\mu}[n]$ with $\widehat{\mu}^{2}[n]\geq(1+\frac{v^{4}[n]}{4v_{0}^{4}})^{\frac{1}{2}}-\frac{v^{2}[n]}{2v_{0}^{2}}$ to obtain the upper bound of $E_{\mathrm{ua}}^{\mathrm{fly}}$, which is expressed as
@@
$g(\widehat{\mu},v[n])=(\widehat{\mu}^{(l)}[n])^{2}+2\mu^{(l)}[n](\widehat{\mu}[n]-\widehat{\mu}^{(l)}[n])$
$+\frac{2}{v_{0}^{2}\delta_{\mathrm{t}}^{2}}(\mathbf{q}_{\mathrm{ua}}^{(l)}[n+1]-\mathbf{q}_{\mathrm{ua}}^{(l)}[n])^{T}(\mathbf{q}_{\mathrm{ua}}[n+1]-\mathbf{q}_{\mathrm{ua}}[n])$
$-\frac{\left\|\mathbf{q}_{\mathrm{ua}}^{(l)}[n+1]-\mathbf{q}_{\mathrm{ua}}^{(l)}[n]\right\|^{2}}{v_{0}^{2}\delta_{\mathrm{t}}^{2}}\geq\frac{1}{\widehat{\mu}^{2}[n]}.$ (47)
As a result, the optimization problem (39) in the $(l+1)$-th iteration of the Dinkelbach’s algorithm can be transformed as
$\max_{\boldsymbol{\Xi}}~{}~{}~{}\mathbf{L}(\boldsymbol{\Xi})-\psi^{(l)}\widehat{E}_{\mathrm{tol}}(\boldsymbol{\Xi}),$